40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10257     Entry time: Tue Jul 22 23:10:12 2014
Author: Akhil 
Type: Update 
Category: General 
Subject: Weekly Update 

 Work Done:

  • Created a Channel Access Server on the Raspberry Pi  to write data from the FC into EPICS Channel.
  • Completed characterization and noise estimation of the FC counter with improved timing.
  • Started installation of FC inside the 40m.

Plans for this Week:

  • Testing how well the FC can replace the spectrum analyzer which is in the control room. For this I have asked Steve to order  an RF adder/combiner to see how frequency counter responds to two RF signals at different frequencies(much like the RF signal fed to the spectrum analyzer) .
  • Complete the installation of FC insode the 40m and start initial testing.
  • Characterization of the Temperature Actuator and initial PID loop design.

Inside the 40m Lab:

  • I will have to go inside the 40m lab this week for routing the RF mon cables to the FC box(in detail:http://nodus.ligo.caltech.edu:8080/40m/10163) .
  • Also to setup for characterization of the temperature actuator, I will be required to go inside the lab in this week.
ELOG V3.1.3-